Find I (the rate period) and n (the number of periods) for the following annuity.

Monthly deposits of $305 are made for 9 years into an annuity that pays 6% compounded monthly.
i = ?
Answer #1

Rate Per period(i) = Rate Per Year /12 = 0.06/12 =0.005 = 0.50%

Number of Periods (n)= Number of Years*12 = 9*12 = 108